OpenStudy (anonymous):

5 to the power of negative 8 times 5 ttpo 4

OpenStudy (anonymous):

\[if you meant: 5^{-8}*5^{4}\] Then the answer is \[5^{-4}\]

OpenStudy (anonymous):

what is that in a fraction?

OpenStudy (anonymous):

I.e you add the exponents ( -8 + 4 =-4)

OpenStudy (anonymous):

\[1/5^{4}\]

OpenStudy (anonymous):

in a fraction it would be written as 

OpenStudy (anonymous):

\\[5^{-8} * 5^{4} = 1/5^{8} + 5^{4} = 5^{(-8+4)} = 5^{-4} = 1 \div5^{4} \][1\div5^{4}\]